OpenStudy (anonymous):

A boy can accelerate at 1.00 m/s2 over a short distance. If the boy were to take an energy pill and suddenly have the ability to accelerate at 5.6 m/s2 then how would his new energy pill force compare his earlier force? If the boy's earlier force was 45N, what is the size of his energy pill force?

OpenStudy (anonymous):

Refer here: http://openstudy.com/study#/updates/4f690653e4b0f81dfbb5d392

OpenStudy (anonymous):

first calculate mass from given initial acceleration by using newtons 2nd law by putting a=1.00m/s^2 & f=45,now again use 2nd law of motion,put the obtained mass in it and a=5.6m/s^2... u will get ur answer.

OpenStudy (mos1635):

F1=m*a1 F2=m*a2 divide those two an you have F1/F2=a1/a2 F1/F2=1/5.6=5/28 F2=F1*(28/5) F2=5.6 * F1 (that is the ans to 1st question(how would his new energy pill force compare his earlier force?)) then if F1=45N then F2=5.6*45 F2=252N

OpenStudy (anonymous):

@mos1635 's solution is almost right, but his F2=252 N is the total force, so you must subtract 45 N from it, therefore 207 N.